4. It is important that face masks used by firefighters be able to withstand high tem- peratures because firefighters commonly work in temperatures of 200-500 degrees. pop out one t of masK, 55 were found to have their lenses at 325 degrees. Construct and interpret a 90% confidence interval for the true proportion of masks of this type whose lenses would pop out at 325 degrees.
Answer #1

Here

sample proportion of people who found ti have their lenses pop out at 325 degrees 18 p=-= 0.327273

and sample size n = 55

a 90% confidence interval for the proportion of people who found ti have their lenses pop out at 325 degrees

=(\hat{p}-z_{0.05}*\sqrt{\frac{(\hat{p}*(1-(\hat{p})}{n}}, \hat{p}+z_{0.05}*\sqrt{\frac{(\hat{p}*(1-(\hat{p})}{n}})

=(0.327273-z_{0.05}*\sqrt{\frac{(0.327273*(1-(0.327273)}{55}}, 0.327273+z_{0.05}*\sqrt{\frac{(0.327273*(1-(0.327273)}{55}})

=(0.327273-1.95996*\sqrt{\frac{(0.327273*(1-(0.327273)}{55}}, 0.327273+1.95996*\sqrt{\frac{(0.327273*(1-(0.327273)}{55}})

=(0.203267, 0.451278)